# Towards Fair RAG: On the Impact of Fair Ranking in Retrieval-Augmented Generation

### v2

@@ 6 Discussion and Conclusion
Why do we often see higher system utility in fairer RAG models? Although there is a general trend of a fairness-utility tradeoff, we observe that certain levels of fairness can actually improve the utility of a baseline RAG model. Recent line of research have uncovered relevant findings: 1) generators are not robust to changes in the position of useful information [30]; 2) items with high retrieval scores often include distracting content that can reduce the system-effectiveness [9, 41]; and 3) introducing some random documents can significantly boost the utility of RAG [9].
@@
Machine-user browsing model. Developing more sophisticated machine-user browsing models will result in more consistent and accurate evaluations of item-side fairness in RAG models, as the exposure of each item is influenced by the attention allocated by the machine-user. Initial research can draw inspiration from Liu et al. [30], who found that LLMs tend to allocate more attention to the beginning and end of a context, with less focus on the middle. This line of inquiry aligns with the broader effort to create search engines tailored for machine-users [44], specifically focusing on fairer search engines in this context. It should involve studying how LMs attend to each retrieved result within a context, analogous to how traditional search engine research models human browsing behavior [11].
@@
Measurement of string utility. In line with the recent call for evaluating various valid output strings [58], we recognize the need for a similar approach to better measure system utility across different rankings given. Recall that our experiments were designed to provide the generator with different rankings for the same query, leading to varied outputs. This approach is motivated by the idea that items not appearing in the top positions of deterministic rankings may still hold value and should be fairly considered by the system. In this context, the diverse outputs generated from different rankings may still be valid. However, we currently rely on a single target output string for comparison with predictions. Future work could focus on calculating the utility of diffuse predictions, enabling a more nuanced evaluation.
@@
Limitations. We acknowledge that the evaluation cost of fair RAG systems can be high due to repeated sampling and inference steps. However, in production, only a single ranking is sampled, minimizing the impact on system latency (Appendix E). Also, a limitation in our utility labeling is that it considers single items, while multiple items may yield contrasting utility gains. Despite this, the strong correlation between ranking quality and system effectiveness suggests this approach reasonably approximates item-worthiness for evaluating the impact of fair ranking on RAG systems.
@@
Conclusion. This study highlights the impact of fair rankings on both the ranking and generation quality of RAG systems. Through the extensive analysis, we show that fairer RAG systems not only maintain high generation quality but can also outperform traditional RAG models, challenging the notion of a strict tradeoff between fairness and effectiveness. Our findings provide valuable insights for developing responsible and equitable RAG systems and pave the way for future research in fair ranking and retrieval-augmented generation. In future work, we hope to extend this framework to consider graded or missing judgments and exploring the different notions of fairness in RAG systems, ultimately advancing the field of trustworthy RAG systems research.
@@ C.1 Normalization of EE-D
$\frac{1}{n}<\frac{1}{n-m}\approx 0$ (8)
For upper bound, recall that the $\mathbf{\epsilon}$ is computed based on samples of rankings from a stochastic policy. For relevance-independent browsing models (e.g., MU), all rankings $\pi\in S_{n}$ have identical exposure norms ${\left\lVert\mathbf{\epsilon}^{\pi}\right\rVert}_{2}^{2}$, where $\mathbf{\epsilon}^{\pi}$ is the exposure of items for ranking $\pi$ sampled from the stochastic policy. Then,
${\left\lVert\mathbf{\epsilon}\right\rVert}_{2}^{2}$ $={\left\lVert\mathbb{E}_{\pi}[\mathbf{\epsilon}^{\pi}]\right\rVert}_{2}^{2}$ (9)
$\leq\mathbb{E}_{\pi}[{\left\lVert\mathbf{\epsilon}^{\pi}\right\rVert}_{2}^{2}]$ (10)
$=\mathbb{E}_{\pi}[{\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}]={\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}$ (11)
@@
With MU, the upper bound of EE-D, ${\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}$ becomes $\sum_{i=1}^{n}{\text{MU}(i)}^{2}=k$. Therefore, per query $q$, we calculate a normalized EE-D
$\overline{\text{EE-D}}_{q}={\left\lVert\mathbf{\epsilon}\right\rVert}_{2}^{2}/k\quad\in[0,1]$ (12)
C.2 Normalization of EE-R
The ranking quality measure EE-R should be normalized by its true upper and lower bound.
@@ C.2 Normalization of EE-R
The lower bound is achieved when $\mathbf{\epsilon}$ becomes $\mathbf{\epsilon}^{-}$, which is an exposure vector of the worst case ranking $\pi^{-}$ (permutations that rank all non-relevant items above relevant items). Given the assumption made from equation 8 and, $\textrm{C}^{+}$ and $\textrm{C}^{-}$, which are set of indices of relevant and non-relevant items, respectively,
$\langle\mathbf{\epsilon}^{-},\mathbf{\epsilon}^{*}\rangle$ $=\sum^{n}_{i=1}\mathbf{\epsilon}^{-}_{i}\mathbf{\epsilon}^{*}_{i}$ (13)
$\approx\sum_{i\in\textrm{C}^{+}}0\mathbf{\epsilon}^{*}_{i}+\sum_{i\in\textrm{C}^{-}}\mathbf{\epsilon}^{-}_{i}0=0$ (from 8) (14)
@@
$\langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle$ $=\sum^{n}_{i=1}\mathbf{\epsilon}_{i}\mathbf{\epsilon}^{*}_{i}=\sum_{i\in\textrm{C}^{+}}\mathbf{\epsilon}_{i}\mathbf{\epsilon}^{*}_{i}$ (from 8) (15)
$=\mathbf{\epsilon}^{*}_{d}\sum_{i\in\textrm{C}^{+}}\mathbf{\epsilon}_{i}$ (equal exposure principle) (16)
$=\mathbf{\epsilon}^{*}_{d}\sum_{i\in\textrm{C}^{+}}\sum_{\pi^{\prime}\in S^{*}_{n}}w_{\pi^{\prime}}\mathbf{\epsilon}^{\prime}_{i}$ (17)
$=\mathbf{\epsilon}^{*}_{d}\sum_{\pi^{\prime}\in S^{*}_{n}}w_{\pi^{\prime}}\sum_{i\in\textrm{C}^{+}}\mathbf{\epsilon}^{\prime}_{i}$ (18)
$\leq\mathbf{\epsilon}^{*}_{d}\sum_{\pi^{\prime}\in S^{*}_{n}}w_{\pi^{\prime}}(m\mathbf{\epsilon}^{*}_{d})$ (since $\mathbf{\epsilon}^{\prime}$ is optimal) (19)
$=\mathbf{\epsilon}^{*}_{d}(m\mathbf{\epsilon}^{*}_{d})$ (since $\sum_{\pi\in S^{*}_{n}}w_{\pi^{\prime}}=1$) (20)
$=\sum_{i\in\textrm{C}^{+}}\mathbf{\epsilon}^{*}_{i}\mathbf{\epsilon}^{*}_{i}={\left\lVert\mathbf{\epsilon}^{*}\right\rVert}_{2}^{2}$ (21)
@@
$\sum_{i=\textrm{C}^{+}}1^{2}+\sum_{i=\textrm{C}^{-}}(\frac{k-m}{n-m})^{2}=m+\frac{(k-m)^{2}}{n-m}$ (22)
@@
$\sum_{i=\textrm{C}^{+}}(\frac{k}{m})^{2}+\sum_{i=\textrm{C}^{-}}0^{2}=\frac{k^{2}}{m}$ (23)
Therefore, depending on $m$ and $k$, per query $q$, we calculate a normalized EE-R
$\overline{\text{EE-R}}_{q}=\begin{cases}\langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle/(m+\frac{(k-m)^{2}}{n-m})&\text{($m\leq k$)}\\ \langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle/(\frac{k^{2}}{m})&\text{($m>k$)}\end{cases}\in[0,1]$ (24)
C.3 Normalization of EU
Theoretically obtaining a true upper bound of a utility of a RAG model is challenging. Therefore, we approximate the true upper bound by the maximum of the empirically obtained utilities given a fixed RAG model $\langle\mathcal{S},\mathcal{G}\rangle$ with a stochastic ranking sampler.
@@ C.3 Normalization of EU
$u_{max}=max\left(\{u_{\pi}\}_{\pi\in\sigma_{1}}\cup\{u_{\pi}\}_{\pi\in\sigma_{2}}\cup\{u_{\pi}\}_{\pi\in\sigma_{4}}\cup\{u_{\pi}\}_{\pi\in\sigma_{8}}\cup\{u_{\pi}\}_{\pi\in\sigma^{*}}\right)$ (25)
@@
$\overline{\text{EU}}_{q}=\frac{\text{EU}_{q}}{u_{max}}\quad\in[0,1]$ (26)
Normalization of EU is done to get the percentage of closeness to the optimal utility as all the utility values are scaled relative to the maximum value. In other words, the normalized EU value indicates how close the EU is to the maximum utility that the RAG system can get to.
@@
Appendix D Effect of $\alpha$ on Metrics
(a) $\alpha$ Vs. $\overline{\text{EE-D}}$
(b) $\alpha$ Vs. $\overline{\text{EE-R}}$
(c) $\alpha$ Vs. $\overline{\text{EU}}$
Figure 6: Effect of a fairness control parameter $\alpha$ on the (a) disparity of rankings, (b) ranking quality, and (c) generation quality of RAG. The left column is the results on the LaMP task 1 and the right column is on the LaMP task 4, each corresponding to a classification and text generation task, respectively. RAG model is configured with Contriever and Flan-T5-Base for all six figures. Each data point represents the value of the associated metric for one query.
When the three types of plots are observed together, we can infer some interesting observations. In general, we see positive relationships between increasing $\alpha$ and both ranking quality and utility. This implies that we can generally expect a tradeoff between both fairness and ranking quality, as well as between fairness and utility.
However, we can expect some edge cases. For instance, in LaMP-1 (left column of the figure), difference in $\overline{\text{EU}}$ when $\alpha=4$ and $\alpha=8$ is not large, and we can see that even the lower quartile of the utility is increased when $\alpha$ is set to 4 (Figure 6(c)). This can possibly mean that in LaMP-1, the RAG model can maintain considerable utility when the disparity is roughly in the range of [0.6, 0.8]. Similar observation can be made for the LaMP-4 (right column of the figure), except that the $\overline{\text{EE-R}}$ is higher when $\alpha=4$ than when $\alpha=8$ (Figure 6(b)). This indicates that the deterministic retriever does not always provide the maximum ranking quality, and the retriever can sometimes provide higher ranking quality by being more fair, ultimately maintaining considerable or higher utility (similar $\overline{\text{EU}}$ when $\alpha$=4 and when $\alpha$=8) at the same time.
With this preliminary observations in mind, we could delve deeper into the relationships between fairness, ranking quality, and utility, by visualizing and quantifying the combined results from all the runs ($\alpha=1,2,4,8$).
Appendix E Integrating Fair Rankings into a RAG System in a Production Environment
Figure 7: Designing a RAG system that incorporates stochastic fair rankings involves using stochastic sampling, where $N$ can be set to 1 to provide a single ranking to the language model. This can result in a different ranking compared to the one in Figure 1, exposing $d_{1}$ and $d_{3}$ to the language model. This paper is focusing on the evaluation of this system as depicted in the Figure 2. The query and prompt generators are omitted in the figure for brevity.
Appendix F Fairness Vs. Ranking Quality
F.1 Visualization of EE-D Vs. EE-R of FlanT5Small
Figure 8: LaMP 1
Figure 9: LaMP 2
Figure 10: LaMP 3
Figure 11: LaMP 4
Figure 12: LaMP 5
Figure 13: LaMP 6
Figure 14: LaMP 7
F.2 Visualization of EE-D Vs. EE-R of FlanT5Base
Figure 15: LaMP 1
Figure 16: LaMP 2
Figure 17: LaMP 3
Figure 18: LaMP 4
Figure 19: LaMP 5
Figure 20: LaMP 6
Figure 21: LaMP 7
F.3 Visualization of EE-D Vs. EE-R of FlanT5XXL
Figure 22: LaMP 1
Figure 23: LaMP 2
Figure 24: LaMP 3
Figure 25: LaMP 4
Figure 26: LaMP 5
Figure 27: LaMP 6
Figure 28: LaMP 7
F.4 Quantification of Fairness-Ranking Quality Tradeoff
[…]
Task FlanT5Small FlanT5Base FlanT5XXL
BM25 SPLADE Contriever BM25 SPLADE Contriever BM25 SPLADE Contriever
LaMP-1 0.2113 / 0.2546 0.2396 / 0.3147 0.2358 / 0.2102 0.1546 / 0.3574 0.2252 / 0.4043 0.1834 / 0.3594 0.1358 / 0.3390 0.2409 / 0.4130 0.1945 / 0.4009
LaMP-2 0.1599 / 0.3079 0.1863 / 0.3330 0.2072 / 0.3372 0.1665 / 0.3699 0.1899 / 0.3918 0.2269 / 0.4098 0.1651 / 0.4161 0.1834 / 0.4222 0.2029 / 0.4328
LaMP-3 0.0309 / 0.3820 0.0353 / 0.3970 0.0501 / 0.3911 0.0403 / 0.4731 0.0290 / 0.4760 0.0555 / 0.4745 0.0798 / 0.3169 0.1010 / 0.3483 0.1338 / 0.3502
LaMP-4 0.1271 / 0.3425 0.1702 / 0.3741 0.1858 / 0.3793 0.1054 / 0.3812 0.1715 / 0.4236 0.1837 / 0.4303 0.0882 / 0.3669 0.1286 / 0.3973 0.1363 / 0.4035
LaMP-5 0.1058 / 0.3396 0.1014 / 0.3386 0.1031 / 0.3356 0.0946 / 0.3477 0.1144 / 0.3597 0.1157 / 0.3533 0.2072 / 0.4026 0.2071 / 0.4100 0.2086 / 0.4040
LaMP-6 0.1194 / 0.4507 0.1036 / 0.4477 0.1090 / 0.4491 0.1456 / 0.4880 0.1264 / 0.4967 0.1438 / 0.4942 0.1615 / 0.5685 0.1554 / 0.5813 0.1467 / 0.5747
LaMP-7 0.0962 / 0.4984 0.1178 / 0.4976 0.0979 / 0.4894 0.0744 / 0.3811 0.1173 / 0.3908 0.0924 / 0.3926 0.1174 / 0.4800 0.0216 / 0.4784 0.0627 / 0.4885
Table 3: Values on the left are the gradient of a linear line fit to the data points where x-axis is $\overline{\text{EE-D}}$ and y-axis is $\overline{\text{EE-R}}$. Higher the value, stronger the tradeoff between fairness and ranking quality. Values on the right are the DR-AUC on the disparity-ranking quality ($\overline{\text{EE-D}}$ Vs. $\overline{\text{EE-R}}$) curve. Higher the value, stronger the ranking quality, given consistent tradeoff between fairness and relevance.
Appendix G Ranking Quality Vs. Utility of RAG Models
G.1 Visualization of the EE-R Vs. EU
Figure 29: LaMP 1
Figure 30: LaMP 2
Figure 31: LaMP 3
Figure 32: LaMP 4
Figure 33: LaMP 5
Figure 34: LaMP 6
Figure 35: LaMP 7
G.2 Quantification of the Relationship Between Ranking Quality and Utility
[…]
Task BM25 SPLADE Contriever
FlanT5Small FlanT5Base FlanT5XXL FlanT5Small FlanT5Base FlanT5XXL FlanT5Small FlanT5Base FlanT5XXL
LaMP-1 0.3813 / 0.3705 0.8250 / 0.7491 0.7612 / 0.7044 0.1434 / 0.3760 0.7783 / 0.7382 0.7454 / 0.7330 0.1864 / 0.3514 0.7661 / 0.7304 0.7089 / 0.7166
LaMP-2 0.5543 / 0.4130 0.2858 / 0.3218 0.2646 / 0.4495 0.5539 / 0.3941 0.3310 / 0.3407 0.1979 / 0.4787 0.6908 / 0.4567 0.2325 / 0.3600 0.3342 / 0.4772
LaMP-3 0.2500 / 0.9153 0.2134 / 0.9204 0.1633 / 0.8974 0.2544 / 0.9124 0.2061 / 0.9203 0.1737 / 0.9009 0.2538 / 0.9125 0.2025 / 0.9205 0.1568 / 0.8977
LaMP-4 0.2708 / 0.2711 0.2881 / 0.2679 0.1947 / 0.3720 0.2638 / 0.2817 0.3272 / 0.2942 0.2193 / 0.3961 0.2264 / 0.2751 0.2986 / 0.2887 0.2190 / 0.3926
LaMP-5 -0.0125 / 0.3723 -0.0307 / 0.4888 0.1673 / 0.5591 -0.0035 / 0.3849 -0.0160 / 0.4773 0.2044 / 0.5570 0.0090 / 0.3823 -0.0474 / 0.4746 0.1590 / 0.5387
LaMP-6 0.2575 / 0.3883 0.2690 / 0.3953 0.3078 / 0.5445 0.2522 / 0.3737 0.2735 / 0.3960 0.2780 / 0.5372 0.2790 / 0.3783 0.2620 / 0.3891 0.3053 / 0.5356
LaMP-7 0.3229 / 0.5240 0.0889 / 0.6391 0.0157 / 0.6082 0.2781 / 0.5177 0.2029 / 0.6435 -0.0623 / 0.5601 0.2713 / 0.5101 0.0788 / 0.6266 -0.0466 / 0.5723
Table 4: Values on the left are the gradient of a linear line fit to the data points where x-axis is $\overline{\text{EE-R}}$ and y-axis is $\overline{\text{EU}}$. Higher the value, stronger the tradeoff between retrieval quality and generation quality. Values on the right are the RU-AUC on the ranking quality-utility ($\overline{\text{EE-R}}$s. $\overline{\text{EU}}$) curve. Higher the value, stronger the general end-performance of a RAG model when every level of relevance is considered.
Appendix H Item-Fairness Vs. Utility of RAG Models
H.1 Visualization of EE-D Vs. EU
Figure 36: LaMP 1
Figure 37: LaMP 2
Figure 38: LaMP 3
Figure 39: LaMP 4
Figure 40: LaMP 5
Figure 41: LaMP 6
Figure 42: LaMP 7
H.2 Quantification of Fairness-Utility Tradeoff
[…]
Task BM25 SPLADE Contriever
FlanT5Small FlanT5Base FlanT5XXL FlanT5Small FlanT5Base FlanT5XXL FlanT5Small FlanT5Base FlanT5XXL
LaMP-1 0.0693 / 0.1994 0.2254 / 0.6110 0.1673 / 0.5688 0.0851 / 0.2519 0.1998 / 0.6644 0.2385 / 0.6362 0.0456 / 0.1866 0.2016 / 0.6248 0.2413 / 0.6061
LaMP-2 0.1295 / 0.2561 0.0740 / 0.2717 0.0600 / 0.4294 0.1119 / 0.2722 0.0637 / 0.2786 0.0626 / 0.4582 0.1683 / 0.3082 0.0870 / 0.3139 0.1002 / 0.4338
LaMP-3 0.0259 / 0.8738 -0.0012 / 0.8964 0.0101 / 0.8706 0.0118 / 0.8744 0.0014 / 0.9003 0.0280 / 0.8786 0.0198 / 0.8685 0.0063 / 0.9022 0.0264 / 0.8765
LaMP-4 0.0606 / 0.2178 0.0789 / 0.2282 0.0734 / 0.3622 0.0937 / 0.2408 0.1245 / 0.2594 0.0895 / 0.3802 0.1016 / 0.2429 0.1209 / 0.2595 0.1051 / 0.3784
LaMP-5 0.0293 / 0.3995 0.0445 / 0.4857 0.0363 / 0.5453 0.0437 / 0.3989 0.0266 / 0.4763 0.0320 / 0.5466 0.0533 / 0.3930 0.0300 / 0.4943 0.0432 / 0.5495
LaMP-6 0.1177 / 0.3698 0.0753 / 0.3861 0.0521 / 0.5535 0.0806 / 0.3629 0.0623 / 0.3888 0.0561 / 0.5558 0.0823 / 0.3580 0.0597 / 0.3849 0.0616 / 0.5468
LaMP-7 0.0215 / 0.4928 0.0392 / 0.6212 0.0854 / 0.6068 0.0718 / 0.4901 0.0256 / 0.6094 -0.0668 / 0.5714 0.0324 / 0.4839 -0.0151 / 0.6101 -0.0114 / 0.5809
Table 5: Values on the left are the gradient of a linear line fit to the data points where x-axis is $\overline{\text{EE-D}}$ and y-axis is $\overline{\text{EU}}$. Higher the value, stronger the tradeoff between item-fairness and generation quality. Values on the right are the DU-AUC on the disparity-utility ($\overline{\text{EE-D}}$ Vs. $\overline{\text{EU}}$) curve. Higher the value, stronger the general end-performance of a RAG model when every level of fairness is considered.
Appendix I Performance of RAG Models with Varying Fairness Levels
Fairness Intervals
@@ Appendix I Performance of RAG Models with Varying Fairness Levels
Table 6: LaMP-1
Fairness Intervals
@@
Table 7: LaMP-2
Fairness Intervals
@@
Table 8: LaMP-3
Fairness Intervals
@@
Table 9: LaMP-4
Fairness Intervals
@@
Table 10: LaMP-5
Fairness Intervals
@@
Table 11: LaMP-6
Fairness Intervals

### v3

@@ 6.1. Higher System Utility with Fair Rankings
Although there is a general trend of a fairness-utility tradeoff, we observe that certain levels of fairness can actually improve the utility of a baseline RAG model. Recent line of research have uncovered relevant findings: 1) generators are not robust to changes in the position of useful information (Liu et al., 2024); 2) items with high retrieval scores often include distracting content that can reduce the system-effectiveness (Cuconasu et al., 2024; Ru et al., 2024); and 3) introducing some random documents can significantly boost the utility of RAG (Cuconasu et al., 2024).
@@
6.2. Attribution Rate and Consumption Fairness
RAG practitioners should consider both the expected utility (EU) of the generated text and the expected attributed exposure disparity ($\overline{\text{EAE-D}}$) to ensure a balance between utility and fairness in the system’s output. Achieving this balance requires careful adjustment of retrieval fairness to find an interval where the generated content remains both useful and fair in its exposure of sources. One important metric to monitor in this process is the attribution rate, which reflects the proportion of retrieved items that are actually cited or used in the final output.
Monitoring expected attribution rate (EAR) provides valuable insights for understanding how fairness manifests in the final output, offering an essential complement to $\overline{\text{EAE-D}}$. For example, even if a ranking appears fair at the retrieval stage, a low attribution rate indicates that the generator ultimately utilizes only a small subset of the retrieved items, which can amplify disparities among the attributed sources. In such cases, fairness concerns at the generation stage may outweigh those at the retrieval stage. Conversely, when the generator attributes a larger portion of the retrieved set, disparities among the attributed sources may become less pronounced, as the system distributes attention more evenly across a broader set of items.
By comparing expected attribution rate with consumption fairness, practitioners can better disentangle the root causes of imbalances in the final output. Specifically, they can determine whether disparities stem from the retrieval stage, such as an inherently skewed ranking, or from the generator’s behavior in selectively focusing on only a few documents. This comparison can help diagnose fairness issues more effectively, enabling practitioners to refine both retrieval and generation components of a RAG system to achieve more equitable outcomes in the generated content.
6.3. Measuring Consumption Fairness
Ensuring consumption fairness while maintaining high utility of the generated text is a fundamental objective in building a fair RAG system. To achieve this, we measure the exposure of the final attributed items, as these are the items that human users will actually encounter in the system’s output. This approach aligns with the real-world operation of RAG systems, particularly in applications like conversational QA or domain-specific assistants for general information or shopping, where the generated text is accompanied by explicitly cited documents or attributed items. In such cases, measuring the fair exposure of the attributed items is critical, as consumption fairness can have a more direct impact on users compared to retrieval fairness, given that human users only consume what is explicitly surfaced by the system.
A key factor in accurately evaluating exposure-based retrieval fairness is the development of sophisticated machine-user browsing models. The work of Liu et al. (2024) provides valuable insights for designing such advanced browsing models. In our experiments, we employed encoder-decoder models with a short retrieval context to get close to the assumption that equal attention is given to all top-$k$ retrieved items (Liu et al., 2024). However, as the use of long-context becomes prevalent, accurately measuring exposure-based retrieval fairness becomes more challenging due to the generator’s tendency to allocate unequal attention to different items.
Despite these challenges, it is important to highlight that measuring consumption fairness is not dependent on machine-user browsing models. Unlike retrieval fairness, consumption fairness can be measured effectively in long-context models, regardless of how attentions are distributed across items by the generators. Therefore, focusing on consumption fairness ensures that the evaluation remains robust, even in scenarios where exposure patterns vary significantly across different models.
6.4. Measurement of String Utility
@@
In line with the recent call for evaluating various valid output strings (Zhang et al., 2024a), we recognize the need for a similar approach to better measure system utility across different rankings given. Recall that our experiments were designed to provide the generator with different rankings for the same query, leading to varied outputs. This approach is motivated by the idea that items not appearing in the top positions of deterministic rankings may still hold value and should be fairly considered by the system. In this context, the diverse outputs generated from different rankings may still be valid. However, we currently rely on a single target output string for comparison with predictions. Future work could focus on calculating the utility of diffuse predictions, enabling a more nuanced evaluation.
6.5. Limitations
@@
We acknowledge that the evaluation cost of fair RAG systems can be high due to repeated sampling and inference steps. However, in production, only a single ranking is sampled, minimizing the impact on system latency. Also, a limitation in our utility labeling is that it considers single items, while multiple items may yield contrasting utility gains. Despite this, the strong correlation between ranking quality and system effectiveness suggests this approach reasonably approximates item-worthiness for evaluating the impact of fair ranking on RAG systems.
Another limitation lies in the ability to capture the exposure of items beyond those retrieved by the system. Since it is difficult to verify attribution across the entire corpus, our method focuses on measuring the exposure fairness of the attributed sources (e.g., those explicitly cited in the generated output). While this approach does not account for the broader set of potential exposures, it is useful for common RAG applications where the attributed sources are directly displayed to users. In such cases, ensuring fairness among these sources is critical, as they are likely to be the primary content users engage with.
7. Conclusion
@@
This study highlights the impact of fair rankings not only on the ranking and generation quality of RAG systems but also on the equitable attribution of sources in the final output. Through extensive analysis, we demonstrate that fairer RAG models can maintain—and in some cases even surpass—the generation quality of traditional approaches, challenging the assumption of an inherent tradeoff between fairness and effectiveness. Our findings emphasize the importance of fair attribution, showing how improvements at the retrieval stage translate into more equitable exposure of the sources that appear in the generated text. By addressing disparities in both retrieval and attribution, we provide valuable insights for developing responsible and equitable RAG systems.
In future work, we hope to extend this framework to consider graded or missing judgments and exploring the different notions of fairness in RAG systems, ultimately advancing the field of trustworthy RAG systems research.
@@ C.1. Normalization of EE-D
(12) $\frac{1}{n}<\frac{1}{n-m}\approx 0$
For upper bound, recall that the $\mathbf{\epsilon}$ is computed based on samples of rankings from a stochastic policy $\pi$. For relevance-independent browsing models (e.g., MU), all rankings $\pi\in S_{n}$ have identical exposure norms ${\left\lVert\mathbf{\epsilon}^{\pi}\right\rVert}_{2}^{2}$, where $\mathbf{\epsilon}^{\pi}$ is the exposure of items for ranking sampled from $\pi$. Then,
(13) ${\left\lVert\mathbf{\epsilon}\right\rVert}_{2}^{2}$ $={\left\lVert\mathbb{E}_{\pi\sim\pi}[\mathbf{\epsilon}^{\pi}]\right\rVert}_{2}^{2}$
(14) $\leq\mathbb{E}_{\pi\sim\pi}[{\left\lVert\mathbf{\epsilon}^{\pi}\right\rVert}_{2}^{2}]$
(15) $=\mathbb{E}_{\pi\sim\pi}[{\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}]={\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}$
@@
With MU, the upper bound of EE-D, ${\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}$ becomes $\sum_{i=1}^{n}{\text{MU}(i)}^{2}=k$. Therefore, per query $q=\phi_{q}(x)$, we calculate a normalized EE-D
(16) $\overline{\text{EE-D}}={\left\lVert\mathbf{\epsilon}\right\rVert}_{2}^{2}/k\quad\in[0,1]$
The same normalization process can be used for EAE-D, but with attributed items.
C.2. Normalization of EE-R
The relevance measure EE-R should be normalized by its true upper and lower bound.
@@ C.2. Normalization of EE-R
The lower bound is achieved when $\mathbf{\epsilon}$ becomes $\mathbf{\epsilon}^{-}$, which is an exposure vector of the worst case deterministic ranking $\pi^{-}$ (permutations that rank all non-relevant items above relevant items). Given the assumption made from equation 12 and, $\textrm{C}^{+}$ and $\textrm{C}^{-}$, which are set of indices of relevant and non-relevant items, respectively,
(17) $\langle\mathbf{\epsilon}^{-},\mathbf{\epsilon}^{*}\rangle$ $=\sum^{n}_{i=1}\mathbf{\epsilon}^{-}_{i}\mathbf{\epsilon}^{*}_{i}$
(18) $\approx\sum_{i\in\textrm{C}^{+}}0\mathbf{\epsilon}^{*}_{i}+\sum_{i\in\textrm{C}^{-}}\mathbf{\epsilon}^{-}_{i}0=0$ (from 12)
@@
(19) $\langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle$ $=\sum^{n}_{i=1}\mathbf{\epsilon}_{i}\mathbf{\epsilon}^{*}_{i}=\sum_{i\in\textrm{C}^{+}}\mathbf{\epsilon}_{i}\mathbf{\epsilon}^{*}_{i}$ (from 12)
(20) $=\mathbf{\epsilon}^{*}_{d}\sum_{i\in\textrm{C}^{+}}\mathbf{\epsilon}_{i}$ (equal exposure principle)
(21) $=\mathbf{\epsilon}^{*}_{d}\sum_{i\in\textrm{C}^{+}}\sum_{\pi^{\prime}\in S^{*}_{n}}w_{\pi^{\prime}}\mathbf{\epsilon}^{\prime}_{i}$
(22) $=\mathbf{\epsilon}^{*}_{d}\sum_{\pi^{\prime}\in S^{*}_{n}}w_{\pi^{\prime}}\sum_{i\in\textrm{C}^{+}}\mathbf{\epsilon}^{\prime}_{i}$
(23) $\leq\mathbf{\epsilon}^{*}_{d}\sum_{\pi^{\prime}\in S^{*}_{n}}w_{\pi^{\prime}}(m\mathbf{\epsilon}^{*}_{d})$ (since $\mathbf{\epsilon}^{\prime}$ is optimal)
(24) $=\mathbf{\epsilon}^{*}_{d}(m\mathbf{\epsilon}^{*}_{d})$ (since $\sum_{\pi\in S^{*}_{n}}w_{\pi^{\prime}}=1$)
(25) $=\sum_{i\in\textrm{C}^{+}}\mathbf{\epsilon}^{*}_{i}\mathbf{\epsilon}^{*}_{i}={\left\lVert\mathbf{\epsilon}^{*}\right\rVert}_{2}^{2}$
@@
(26) $\sum_{i=\textrm{C}^{+}}1^{2}+\sum_{i=\textrm{C}^{-}}(\frac{k-m}{n-m})^{2}=m+\frac{(k-m)^{2}}{n-m}$
@@
(27) $\sum_{i=\textrm{C}^{+}}(\frac{k}{m})^{2}+\sum_{i=\textrm{C}^{-}}0^{2}=\frac{k^{2}}{m}$
Therefore, depending on $m$ and $k$, per query $q=\phi_{q}(x)$, we calculate a normalized EE-R
(28) $\overline{\text{EE-R}}=\begin{cases}\langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle/(m+\frac{(k-m)^{2}}{n-m})&\text{($m\leq k$)}\\ \langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle/(\frac{k^{2}}{m})&\text{($m>k$)}\end{cases}\in[0,1]$
C.3. Normalization of EU
Theoretically obtaining a true upper bound of a utility of a RAG model is challenging. Therefore, we approximate the true upper bound by the maximum of the empirically obtained utilities given a fixed RAG model $\text{RAG}(\mathcal{R},\mathcal{G},k)$.
@@ C.3. Normalization of EU
(29) $u_{max}=max\left(\{u_{\pi}\}_{\pi\in\sigma_{1}}\cup\{u_{\pi}\}_{\pi\in\sigma_{2}}\cup\{u_{\pi}\}_{\pi\in\sigma_{4}}\cup\{u_{\pi}\}_{\pi\in\sigma_{8}}\cup\{u_{\pi}\}_{\pi\in\sigma^{*}}\right)$
@@
(30) $\overline{\text{EU}}=\frac{\text{EU}}{u_{max}}\quad\in[0,1]$
Normalization of EU is done to get the percentage of closeness to the optimal utility as all the utility values are scaled relative to the maximum value. In other words, the normalized EU value indicates how close the EU is to the maximum utility that the RAG system get get to.
@@
Appendix D Fairer RAG Models Vs. Deterministic RAG Models
Disparity Intervals
@@
Contriever+FlanUL2 (0.620) -0.16 -0.09 -0.05 +0.02 -0.03
Table 4. LaMP-1
Disparity Intervals
@@
Contriever+FlanUL2 (0.455) -0.02 +0.04 +0.09 +0.08 +0.01
Table 5. LaMP-2
Disparity Intervals
@@
Contriever+FlanUL2 (0.680) -0.01 0.00 +0.01 +0.10 +0.04
Table 6. LaMP-3
Disparity Intervals
@@
Contriever+FlanUL2 (0.449) -0.15 +0.01 -0.01 0.00 +0.02
Table 7. LaMP-4
Disparity Intervals
@@
Contriever+FlanUL2 (0.641) -0.07 +0.05 +0.01 -0.03 -0.02
Table 8. LaMP-5
Disparity Intervals
@@
Contriever+FlanUL2 (0.551) -0.01 +0.04 +0.05 +0.05 +0.08
Table 9. LaMP-6
Disparity Intervals